The length of a rectangle is twice the width. the area is 450yd^2. Find the length and the width.
Answer by checkley77(12844)

L=2W
A=LW
REPLACE L WITH 2W & SOLCE FOIR W.
450=2W*W
2W^2=450
W^2=250/2
W^2=225
W=SQRT225
W=15 ANS.